OpenStudy (anonymous):

SHOW ALL OF THE STEP THAT YOU USE TO SOLVE THE PROBLEM 21. Find the distance between points M(6,16) and Z(-1,14) to the nearest tenth.

OpenStudy (anonymous):

Use the distance formula

OpenStudy (anonymous):

\[D=\sqrt{\left( x _{2}-x _{1} \right)^{2}-\left( y _{2}-y _{1} \right)^{2}}\]

OpenStudy (anonymous):

You use the points you were given and use a pair of coordinates as x1 and y1 and the other as x2 and y2

OpenStudy (anonymous):

Once you do that you have d= sqrt((-1-6)^2+(14-16)^2)

OpenStudy (anonymous):

Do the subtraction and get d=sqrt((-7)^2+(-2)^2)

OpenStudy (anonymous):

Then d=sqrt(49+4) Remember when you square a negative number it's positive.

OpenStudy (anonymous):

So do the last step and add those together

OpenStudy (anonymous):

Okay then. \[D=\sqrt{53} units\]

OpenStudy (anonymous):

If you're told to round, you can approximate it to 7.2801 units